# Real Bers Embedding on the Line: Fisher-Rao Linearization, Schwarzian Curvature, and Scattering Coordinates

### v1

@@ 6.3. Miura factorization
The proof is a direct computation. ∎
@@ 6.7. Riccati-Miura characterization and energy identity
6.8. Local Bers coordinates and transition maps
Theorem 6.14 (Local Bers coordinates).
There exists a neighborhood $U$ of $\mathrm{Id}$ in $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and a neighborhood $V$ of 0 in $W^{\infty,1}(\mathbb{R})$ such that
$\beta^{-\infty}\colon U\longrightarrow V$
is a Fréchet diffeomorphism onto its image, with Fréchet smooth inverse. Equivalently, in the linearizing coordinate $u=\log\varphi^{\prime}$ the map
$\mathcal{B}(u):=\beta^{-\infty}(\Phi_{\infty}^{-1}(u))=\frac{1}{2}\left(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\right)$
is a tame Fréchet local diffeomorphism near $u=0$.
Proof.
Work in the global coordinate $u=\log\varphi^{\prime}\in W^{\infty,1}(\mathbb{R})$ via the global chart $\Phi_{\infty}$. Then $\mathcal{B}\colon W^{\infty,1}\to W^{\infty,1}$ is Fréchet smooth and its differential at $u$ is
$D_{u}\mathcal{B}(h)=\frac{1}{2}\bigl(h^{\prime\prime}-u^{\prime}h^{\prime}\bigr)=\frac{1}{2}\,\mathcal{D}_{u}(h).$
By Theorem 6.7, $\mathcal{D}_{u}$ admits a continuous tame right inverse $R_{u}$ on bounded sets, hence $D_{u}\mathcal{B}$ has a continuous tame right inverse $\frac{1}{2}R_{u}$.
It remains to show that $D_{u}\mathcal{B}$ is injective on $W^{\infty,1}(\mathbb{R})$. If $D_{u}\mathcal{B}(h)=0$, then $\mathcal{D}_{u}(h)=h^{\prime\prime}-u^{\prime}h^{\prime}=0$, hence $(\mathrm{e}^{-u}h^{\prime})^{\prime}=0$. Thus $\mathrm{e}^{-u}h^{\prime}=c$ is constant. Since $h\in W^{\infty,1}(\mathbb{R})$ we have $h^{\prime}\in W^{\infty,1}(\mathbb{R})\subset L^{1}(\mathbb{R})$ and $h^{\prime}(\pm\infty)=0$, so $c=0$ and $h^{\prime}\equiv 0$. Because $h\in L^{1}(\mathbb{R})$, this forces $h\equiv 0$.
Therefore $D_{u}\mathcal{B}$ is a tame linear isomorphism between the model Fréchet spaces at each $u$ in a sufficiently small neighborhood of 0. Applying the Nash-Moser inverse function theorem in tame Fréchet spaces yields a neighborhood of 0 on which $\mathcal{B}$ is a tame local diffeomorphism, and transporting back by $\Phi_{\infty}$ gives the claimed local diffeomorphism property of $\beta^{-\infty}$ near $\mathrm{Id}$. ∎
Corollary 6.15 (Quantitative stability on the Bers chart).
For each $k\geq 0$ there exist neighborhoods $U_{k}\subset\operatorname{Diff}^{-\infty}(\mathbb{R})$ of $\mathrm{Id}$ and $V_{k}\subset W^{k,1}(\mathbb{R})$ of 0 and a constant $C_{k}>0$ such that if $\varphi,\psi\in U_{k}$ and $q_{\varphi}=\beta^{-\infty}(\varphi)$, $q_{\psi}=\beta^{-\infty}(\psi)\in V_{k}$, then
$\|\log\varphi^{\prime}-\log\psi^{\prime}\|_{W^{k+2,1}}\leq C_{k}\,\|q_{\varphi}-q_{\psi}\|_{W^{k,1}}.$
In particular, $\|\varphi^{\prime}-\psi^{\prime}\|_{W^{k+2,1}}\leq C_{k}^{\prime}\|q_{\varphi}-q_{\psi}\|_{W^{k,1}}$ and $\|\varphi-\psi\|_{W^{k+3,1}}\leq C_{k}^{\prime\prime}\|q_{\varphi}-q_{\psi}\|_{W^{k,1}}$.
Proof.
This is an immediate consequence of Theorem 6.14 together with the tame estimates in Theorem 6.7 for the right inverse of the linearization and the smoothness of $u\mapsto\Phi_{\infty}^{-1}(u)$ on bounded sets. ∎
Proposition 6.16 (Transition maps between Bers charts).
Fix $\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$. Define the Bers chart at $\psi$ by
$\beta^{-\infty}_{\psi}(\varphi):=\mathcal{A}_{\psi^{-1}}\bigl(\beta^{-\infty}(\varphi)\bigr).$
Then $\beta^{-\infty}_{\psi}(\psi)=0$ and, on neighborhoods where these charts are defined, the transition map between charts at $\psi$ and $\chi$ is
$\beta^{-\infty}_{\chi}(\varphi)=\mathcal{A}_{\chi^{-1}\circ\psi}\bigl(\beta^{-\infty}_{\psi}(\varphi)\bigr).$
Proof.
Since $\beta^{-\infty}(\psi)=\mathcal{A}_{\psi}0$ and $\mathcal{A}_{\psi^{-1}}\circ\mathcal{A}_{\psi}=\mathcal{A}_{\psi\circ\psi^{-1}}=\mathcal{A}_{\mathrm{Id}}$, we have $\beta_{\psi}^{-\infty}(\psi)=0$. For the transition, use $\mathcal{A}_{\chi^{-1}}=\mathcal{A}_{\chi^{-1}\circ\psi}\circ\mathcal{A}_{\psi^{-1}}$ from Proposition 6.5(i). ∎
6.9. Spectral transport and corrected Liouville conjugation
The next result replaces the incorrect operator-level statement that omits the $(\varphi^{\prime})^{-2}$ prefactor and uses the wrong half-density weight. The correct transport is naturally formulated using the Liouville map (54).
Theorem 6.17 (Spectral transport).
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $q=\beta^{-\infty}(\varphi)=Q(\varphi)$. Define $\mathcal{L}_{\varphi}$ by (54). Then, on $C_{c}^{\infty}(\mathbb{R})$, one has the exact conjugation identity
$\mathcal{L}_{\varphi}\,\partial_{y}^{2}\,\mathcal{L}_{\varphi}^{-1}=(\varphi^{\prime})^{-2}\,\bigl(\partial_{x}^{2}+q(x)\bigr)=(\varphi^{\prime})^{-2}\,L_{q}.$ (64)
Equivalently,
$L_{q}\,\mathcal{L}_{\varphi}=(\varphi^{\prime})^{2}\,\mathcal{L}_{\varphi}\,\partial_{y}^{2}.$ (65)
Moreover, $\mathcal{L}_{\varphi}$ extends to a unitary map
$\mathcal{L}_{\varphi}\colon L^{2}(\mathbb{R},dy)\longrightarrow L^{2}(\mathbb{R},\varphi^{\prime}(x)^{2}\,dx),$
and therefore the self-adjoint realization of $(\varphi^{\prime})^{-2}L_{q}$ on $L^{2}(\mathbb{R},\varphi^{\prime 2}dx)$ is unitarily equivalent to the free Laplacian $\partial_{y}^{2}$ on $L^{2}(\mathbb{R},dy)$.
In particular, for each $\xi\in\mathbb{R}$ the distorted plane waves
$\psi_{\xi}^{\pm}(x):=(\varphi^{\prime}(x))^{-1/2}\,\mathrm{e}^{\pm i\xi\,\varphi(x)}$ (66)
satisfy the exact eigenvalue equation
$(\varphi^{\prime})^{-2}L_{q}\,\psi_{\xi}^{\pm}=-\xi^{2}\,\psi_{\xi}^{\pm},\qquad\text{equivalently}\qquad L_{q}\,\psi_{\xi}^{\pm}=-\xi^{2}\,(\varphi^{\prime})^{2}\,\psi_{\xi}^{\pm}.$
Proof.
Using Proposition 6.6 we write $L_{q}=(\partial_{x}-\frac{u^{\prime}}{2})(\partial_{x}+\frac{u^{\prime}}{2})$ with $u=\log\varphi^{\prime}$. Let $f\in C_{c}^{\infty}(\mathbb{R})$ and set $g=\mathcal{L}_{\varphi}f=\mathrm{e}^{-u/2}(f\circ\varphi)$. A direct computation gives
$\left(\partial_{x}+\frac{u^{\prime}}{2}\right)g=\left(\partial_{x}+\frac{u^{\prime}}{2}\right)\left(\mathrm{e}^{-u/2}f(\varphi(x))\right)=\mathrm{e}^{-u/2}\,f^{\prime}(\varphi(x))\,\varphi^{\prime}(x)=\mathrm{e}^{u/2}\,(f^{\prime}\circ\varphi).$
Applying the second factor yields
$L_{q}g=\left(\partial_{x}-\frac{u^{\prime}}{2}\right)\Bigl(\mathrm{e}^{u/2}(f^{\prime}\circ\varphi)\Bigr)=\mathrm{e}^{u/2}\,\varphi^{\prime}\,(f^{\prime\prime}\circ\varphi)=(\varphi^{\prime})^{3/2}\,(f^{\prime\prime}\circ\varphi).$
Since $\mathcal{L}_{\varphi}(\partial_{y}^{2}f)=(\varphi^{\prime})^{-1/2}(f^{\prime\prime}\circ\varphi)$, this is exactly
$L_{q}\,\mathcal{L}_{\varphi}f=(\varphi^{\prime})^{2}\,\mathcal{L}_{\varphi}(\partial_{y}^{2}f),$
which is (65). Conjugation form (64) follows by right-multiplying by $\mathcal{L}_{\varphi}^{-1}$.
For unitarity, compute for $f\in C_{c}^{\infty}(\mathbb{R})$:
$\|\mathcal{L}_{\varphi}f\|_{L^{2}(\varphi^{\prime 2}dx)}^{2}=\int_{\mathbb{R}}\left|(\varphi^{\prime})^{-1/2}f(\varphi(x))\right|^{2}\,\varphi^{\prime}(x)^{2}\,dx=\int_{\mathbb{R}}|f(\varphi(x))|^{2}\,\varphi^{\prime}(x)\,dx=\int_{\mathbb{R}}|f(y)|^{2}\,dy,$
by the change of variables $y=\varphi(x)$. The eigenfunction statement (66) is obtained by taking $f(y)=\mathrm{e}^{\pm i\xi y}$. ∎
Corollary 6.18 (Trivial scattering for the transported Laplacian).
Let $q=\beta^{-\infty}(\varphi)$. The operator $(\varphi^{\prime})^{-2}L_{q}$ on $L^{2}(\mathbb{R},\varphi^{\prime 2}dx)$ is unitarily equivalent to the free Laplacian on $L^{2}(\mathbb{R},dy)$ and therefore has purely absolutely continuous spectrum $(-\infty,0]$. In the distorted Fourier representation determined by (66), the operator is diagonal with no mode mixing which is equivalent to a trivial monodromy.
Proof.
This is immediate from Theorem 6.17 and the spectral theorem for $\partial_{y}^{2}$ on $L^{2}(\mathbb{R},dy)$. ∎
Lemma 6.19 (Unitary Koopman operator on half-densities).
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, the operator
$(\mathcal{K}_{\varphi}f)(x):=\sqrt{\varphi^{\prime}(x)}\,f(\varphi(x))$
is unitary on $L^{2}(\mathbb{R},dx)$. Moreover $\mathcal{L}_{\varphi}=M_{\varphi^{\prime-1}}\mathcal{K}_{\varphi}$, where $M_{\varphi^{\prime-1}}$ denotes multiplication by $\varphi^{\prime-1}$.
Proof.
Unitarity is the change of variables $y=\varphi(x)$:
$\|\mathcal{K}_{\varphi}f\|_{L^{2}(dx)}^{2}=\int_{\mathbb{R}}\varphi^{\prime}(x)|f(\varphi(x))|^{2}\,dx=\int_{\mathbb{R}}|f(y)|^{2}\,dy.$
The factorization $\mathcal{L}_{\varphi}=M_{\varphi^{\prime-1}}\mathcal{K}_{\varphi}$ is immediate from definitions. ∎
@@ 7.4. The scattering transform as a nonlinear Fourier map
For injectivity, let $\varphi_{1},\varphi_{2}\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $q_{j}=\beta^{-\infty}(\varphi_{j})$. Assume $\mathcal{S}(\varphi_{1})=\mathcal{S}(\varphi_{2})$, so $R_{q_{1}}(k)=R_{q_{2}}(k)$ for all $k\in\mathbb{R}\setminus\{0\}$. By Lemma 7.3, the operators $H_{q_{j}}$ have no negative spectrum, hence their scattering data have no discrete part. The Gel’fand-Levitan-Marchenko reconstruction for short-range one-dimensional Schrödinger operators then implies $q_{1}=q_{2}$ from equality of the reflection coefficients in the no-bound-state case, see [10, Theorem 1]. The injectivity of $\beta^{-\infty}$ modulo $\operatorname{Aff}(\mathbb{R})$ from Theorem 6.2 then gives $\varphi_{1}]=[\varphi_{2}]$ in $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.

### v2

@@ 5.3. Miura factorization
Set $u=\log(\varphi^{\prime})$, so $\varphi^{\prime}=\mathrm{e}^{u}$. By Theorem 5.2,
$q:=\beta^{-\infty}(\varphi)=\frac{1}{2}S(\varphi)=\frac{1}{2}\Bigl(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\Bigr)=\frac{u^{\prime\prime}}{2}-\frac{(u^{\prime})^{2}}{4}.$
Expanding,
$\Bigl(\partial_{x}-\frac{u^{\prime}}{2}\Bigr)\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)=\partial_{x}^{2}+\frac{u^{\prime\prime}}{2}-\frac{(u^{\prime})^{2}}{4}=\partial_{x}^{2}+q=L_{\beta^{-\infty}(\varphi)}.$
For $y_{2}=\mathrm{e}^{-u/2}$ one computes
$\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)y_{2}=\Bigl(-\frac{u^{\prime}}{2}+\frac{u^{\prime}}{2}\Bigr)\mathrm{e}^{-u/2}=0,$
and therefore $L_{\beta^{-\infty}(\varphi)}y_{2}=0$ by the factorization. Since $y_{2}>0$, this is the distinguished positive solution. ∎
@@ 5.7. Riccati-Miura characterization and energy identity
Remark 5.14 (Local coordinates and spectral transport).
Local Bers coordinates centered at arbitrary basepoints can be constructed, with transition maps given by Schwarzian-shifted affine connections. A spectral transport formula conjugates Schrödinger operators along paths in the Bers image. These constructions will appear in a companion paper.
@@ 6.4. The scattering transform as a nonlinear Fourier map
For injectivity, let $\varphi_{1},\varphi_{2}\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $q_{j}=\beta^{-\infty}(\varphi_{j})$. Assume $\mathcal{S}(\varphi_{1})=\mathcal{S}(\varphi_{2})$, so $R_{q_{1}}(k)=R_{q_{2}}(k)$ for all $k\in\mathbb{R}\setminus\{0\}$. By Lemma 6.3, the operators $H_{q_{j}}$ have no negative spectrum, hence their scattering data have no discrete part. The Gel’fand-Levitan-Marchenko reconstruction for short-range one-dimensional Schrödinger operators then implies $q_{1}=q_{2}$ from equality of the reflection coefficients in the no-bound-state case, see [7, Theorem 1]. The injectivity of $\beta^{-\infty}$ modulo $\operatorname{Aff}(\mathbb{R})$ from Theorem 5.2 then gives $\varphi_{1}]=[\varphi_{2}]$ in $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.